7.40 1.60 7.40 7.33/ 7.60 2.81 2.81 7.60 14 7.40 7.60 7.60 HO 7.37 7.37 PPM

What is the multiplicity, integration, number of neighboring Hydrogen for each carbon?
Answer #1

The four aromatic protons at 7.6 ppm have coupled with the adjacent one aromatic proton, so multiplicity = n+1 = 1+1 = 2, i.e. doublet. (multiplicy = 2, integration = 4, number of neighboring hydrogens = 1)

The four aromatic protons at 7.4 ppm have coupled with the adjacent two aromatic protons, so multiplicity = n+1 = 2+1 = 3, i.e. triplet. (multiplicy = 3, integration = 4, number of neighboring hydrogens = 2)

The two alkenyl protons at 7.37 ppm have no coupling partners, so multiplicity = n+1 = 0+1 = 1, i.e. singlet. (multiplicy = 1, integration = 2, number of neighboring hydrogens = 0)

The two aromatic protons at 7.33 ppm have coupled with the adjacent two aromatic protons, so multiplicity = n+1 = 2+1 = 3, i.e. triplet. (multiplicy = 3, integration = 2, number of neighboring hydrogens = 2)

The four methylene protons at 2.81 ppm have coupled with the adjacent two methylene protons, so multiplicity = n+1 = 2+1 = 3, i.e. triplet. (multiplicy = 3, integration = 4, number of neighboring hydrogens = 2)

The two methylene protons at 1.6 ppm have coupled with the adjacent four methylene protons, so multiplicity = n+1 = 4+1 = 5, i.e. quintet. (multiplicy = 5, integration = 2, number of neighboring hydrogens = 4)

Therefore, the total number of protons = 4+4+2+2+4+2 = 18